In a class of 30 students, there are six more boys than girls. What percentage of this class are girls?
Answer by ovie27(16)

Suppose,
x = number of girls
y = number of boys
so, x + y = 30
There are six more boys than girls, so:
y = x + 6
Plug in y into the equation:
x + (x + 6) = 30
2x +6 = 30
2x = 24
x = 12
So, there are 12 girls in the class, and there are 18 boys in the class.
Now the percentage: ( let z equal the unknown percentage)
12/30 = z/100
cross multiply:
1200 = 30z
z = 40 %
So, the percentage of girls in the class is 40 percent.